### v3

@@ 2.2 Unbiased forecasts of counterfactuals
In this section, we characterize the class of DGPs for the counterfactuals. The need to discuss the DGP for counterfactuals arises because of the lack of a control group, which means that we must rely on forecasting counterfactuals from pre-treatment observations.
@@ 2.2.1 Stationary or stochastic trends DGPs
In this section, we consider a class of DGPs such that Assumption 1 is satisfied generally, namely, by any forecast that can be written as a weighted average of pre-treatment outcomes with weights summing to 1. The class of DGPs expresses the counterfactual as the sum of potentially two unobserved stochastic components. This includes a variety of processes, such as stationary and non-stationary (unit root) ARMA processes with individual-specific parameters.
@@
where $y_{it}^{(1)}(0)$ is an unobserved mean-stationary process and $y_{it}^{(2)}(0)=y_{it-1}^{(2)}(0)+u_{it}(0)$ is an unobserved random walk process with innovations satisfying $\mathbb{E}u_{it}(0)=0$, for all $t\geq 2$. Either/both components could be zero.
@@
Assumption 3 does not require both components to be present, which means that it accommodates stationarity as well as non-stationarity processes due to a stochastic trend. The user does not need to take a stance on the component(s). Our method is robust to both. When both components in Assumption 3 are present, the assumption is equivalent to the classical trend-cycle decomposition of macroeconomic time series with stochastic trends (e.g., Nelson and Plosser (1982); Watson (1986)).
@@
This class of DGPs is a plausible assumption for applications where either: 1) the time series of pre-treatment outcomes does not display a trend; 2) there is a trend in pre-treatment outcomes that is plausibly stochastic (not deterministic); 3) there is only one pre-treatment observation so a deterministic trend could never be modelled anyway. We consider processes with deterministic trends in pre-treatment outcomes in the next section.
A key insight of this paper is that a correctly specified model (beyond Assumption 3) is not necessary to obtain unbiased forecasts of the counterfactuals. In fact, as the next result shows, any forecast expressed as a weighted average of pre-treatment observations satisfies the unbiasedness condition.
@@
Note that Theorem 1 shows how to obtain unbiased estimates of the individual (possibly heterogeneous) treatment effects. The requirement in (11) is stronger than the average unbiasedness of Assumption 1. This means that under the assumptions of Theorem 1 one can not only obtain consistent and asymptotically normal estimates of the average effects, but also unbiased estimates of the individual effects.
There are many ways to obtain forecasts that are weighted averages of pre-treatment data - the sample mean being the most obvious example. In this paper we focus on a general class of forecasts obtained via basis function regressions, such as polynomial time trends regressions. This class has the sample mean as a special case.
@@
Consider a sequence of linearly independent functions $\{b_{k}(t)\}_{k=0}^{q_{i}}$, $q_{i}\in\left\{0,1,2,\ldots,\tau-1\right\}$, on the interval $\mathcal{T}_{i}=\left\{\tau-R_{i}+1,\ldots,\tau\right\}$ with $R_{i}\in\left\{q_{i}+1,\ldots,\tau\right\}$, and such that $b_{0}(t)=1$ for all $t$. For example, polynomial time trends set $b_{k}(t)=t^{k}$, with $q_{i}$ the order of the polynomial. For each individual $i$, we forecast the counterfactual via individual-specific regressions of pre-treatment outcomes $\{y_{it}\}_{t\in\mathcal{T}_{i}}$ on the basis functions $\{b_{k}(t)\}_{k=0}^{q_{i}}$:
@@
The definition above makes it clear that for any type of basis function the choice $q_{i}=0$ yields the sample mean of pre-treatment outcomes as the forecast. The following example illustrates how the weighted average representation can arise quite naturally.
Example 1.
Consider $b_{k}(t)=t^{k}$. Set $R_{i}=q_{i}+1$ and $h=1$. In this case, $\widehat{y}_{i\tau+1}^{(q_{i})}(0):=\widehat{y}_{i\tau+1}^{(q_{i},q_{i}+1)}(0)$ can be defined iteratively as:
[…]
This iteration is quite intuitive: the $q_{i}$-forecast $\widehat{y}_{i\tau+1}^{(q_{i})}(0)$ is formed by subtracting the lagged forecast error $\widehat{y}_{i\tau}^{(q_{i}-1)}(0)-y_{i\tau}$ from the $q_{i}-1$ forecast $\widehat{y}_{i\tau+1}^{(q_{i}-1)}(0)$. Explicit formulas for this case are given by
$\widehat{y}_{i\tau+1}^{(0)}(0)$ $=y_{i\tau},$ $\widehat{y}_{i\tau+1}^{(q_{i})}(0)$ $=\sum_{t=\tau-q_{i}}^{\tau}\,w_{it}^{(q_{i})}\,y_{it},$ $w_{it}^{(q_{i})}$ $=(-1)^{(\tau-t)}\,{q_{i}+1\choose\tau-t+1},$
where $\binom{a}{b}=\frac{a!}{b!(a-b)!}$ is the binomial coefficient. [^19] Laderman and Laderman (1982) derive a similar expression in the context of forecasting a time series by polynomial regression using the entire available time series. The weights sum to 1.
We now show that forecasts obtained via basis function regressions satisfy the weighted average requirement of Theorem 1.
@@
For known basis functions $\{b_{k}(t)\}_{k=0}^{q_{i}},\,q_{i}=0,1,\dots,\tau_{i}-1$ that are linearly independent on $\mathcal{T}_{i}$ with $b_{0}\left(t\right)=1$, the forecast in (12) can be written as a weighted average of past outcomes with weights that sum to 1 as in (10) in Theorem 1.
@@ 2.2.2 Deterministic trends DGPs
where $y_{it}^{(1)}(0)$ and $y_{it}^{(2)}(0)$ are as in Assumption 3 and $y_{it}^{(3)}(0)$ is a deterministic time trend $y_{it}^{(3)}(0)=\sum_{k=0}^{q_{0i}}c_{ik}^{(3)}{b}_{k}(t)$ with $c_{i}^{(3)}\in\mathbb{R}^{q_{0i}+1}$ and known basis functions $\{b_{k}(t)\}_{k=0}^{q_{0i}}$, $q_{0i}\in\left\{0,1,2,\ldots,\tau-1\right\}$. We assume that $y_{it}^{(3)}(0)$ is not zero, whereas $y_{it}^{(1)}(0)$ and/or $y_{it}^{(2)}(0)$ could be zero.
@@
where $\widehat{y}_{i\tau+h}^{(q_{i},R_{i})}(0)$ is defined in (12), with $q_{i}\geq q_{0i}$.
@@
While the stochastic components of the counterfactual in Assumption 4 are unobserved and do not have to be present, the deterministic time trend component is assumed to be always present and a function of the same basis functions used to obtain the forecast. This implies that the stochastic component of the DGP does not need to be correctly specified, but, if a deterministic trend component is present, it is correctly specified up to the order (as we discuss in the next remark). This is an unusual result from the perspective of forecasting, where one typically focuses on specifying both stochastic and deterministic parts of a model. [^20] In the next section, we mention how the assumption of correct specification for the deterministic component could in principle be relaxed.
@@
The key requirement of Theorem 2 is that $q_{i}$, the number of basis functions used in estimation, be at least $q_{0i}$, the true number of basis functions. Intuitively, this means that choosing a too small number of basis functions runs the risk of delivering biased forecasts of the counterfactuals. We next discuss the practical implications of this finding.
@@ 2.3 Choice of basis functions and tuning parameters
Our proposed method for forecasting counterfactuals in Definition 1 requires choosing: 1) the type of basis functions $b_{k}(t)$; 2) the number of basis functions $q_{i}$; and 3) the number of pre-treatment periods used for the estimation $R_{i}$. We discuss the tradeoffs that these choices create and offer some practical recommendations for empirical researchers.
According to Theorems 1 and 2, the choice of basis functions only matters when the DGP has a deterministic time trend, in which case the basis functions need to be correctly specified to ensure unbiasedness (up to the order). When the DGP is possibly stationary or has a stochastic trend, the choice of basis functions does not matter for unbiasedness.
@@
Basis functions may be chosen based on the time series properties of pre-treatment outcomes. Polynomial time trends seem to be a natural choice of basis functions for DGPs with deterministic trends. [^21] In applications using difference-in-differences (DiD) methods it is typical to assume the presence of time trends (mostly linear) that are common between control and treatment groups. Our results make it clear that a linear time trend can only be dealt with by either using a control group or, when a control group is not available, by using (at least two) pre-treatment time periods to model the trend (which leads to our polynomial regression) Other basis functions could be used, e.g., periodicity could be captured by Fourier basis functions. Our practical recommendation, and what we focus on henceforth, is to consider polynomial basis functions by letting $b_{k}(t)=t^{k}$ in Definition 1.222222An advantage of polynomial basis functions is that it is in principle possible to relax the assumption that the basis functions are correctly specified, assuming instead that $y_{it}^{(3)}(0)$ in Assumption 4 is a continuous (but unknown) function of time. Since time in our setting is defined on a compact interval and the deterministic trend is a continuous function, $y_{it}^{(3)}(0)$ can be approximated arbitrarily well by a polynomial in time. In fact, by the Weierstrass Approximation Theorem, the approximation error approaches zero as the order of the polynomial goes to infinity. Under additional smoothness assumptions on the deterministic trend, an approximation theorem could then be used (e.g. the Polynomial Approximation Error Theorem) to derive a bound on the approximation error of $y_{it}^{(3)}(0)$ by the polynomial regression. The forecast of $y_{i\tau+h}(0)$ is biased, but we conjecture that it may be possible to do bias correction given an expression for the bias obtained via the approximation theorem.
The order of the basis functions $q_{i}$ used for the estimation only matters if the DGP has a deterministic trend component. For the class of DGPs without a deterministic trend, any $q_{i}$ ensures unbiasedness, and choosing the smallest order, $q_{i}=0$, gives the estimator with the lowest variance. On the other hand, in one of the simulations in Section 4, we find that choosing larger $q_{i}$ could help control the bias resulting from a non-stationary initial condition, so one may want to additionally consider $q_{i}=1,2$ if nonstationary initial conditions are a concern. If the DGP has a deterministic trend component, then $q_{i}$ cannot be smaller than the true order. The true order is unknown and cannot be consistently estimated, so a trade-off emerges where a large $q_{i}$ ensures unbiasedness but comes at the cost of higher variance. Cross-validation methods on pre-treatment data cannot help choose $q_{i}$ if they only target accuracy, because of the necessity to ensure unbiasedness in our context. [^23] It may be possible to devise bias-correction within cross-validation methods to select $q_{i}$, but we leave this endeavour for future work.
Our practical recommendation is to report results for a small range of values for $q_{i}$, e.g., $q_{i}=0,1,2,3$. In addition, plotting the time series of pre-treatment observations can provide some informal guidance on how to choose this range (for example, if the pre-treatment time series displays a time trend that is plausibly deterministic, $q_{i}=0$ can be ruled out; if pre-treatment data appear stationary, any choice of $q_{i}$ is valid).
Regarding the choice of estimation window, $R_{i}$, in our class of DGPs it makes sense to choose $R_{i}$ as large as possible to obtain an estimator with small variance. On the other hand, a short $R_{i}$ can guard against violation of our assumptions due to parameter instability in pre-treatment data. Plotting pre-treatment time series can indicate whether parameter instability is a concern in a given application.
@@ 2.4 Individual treatment timing and limited anticipation
The treatment timing $\tau$ can be individual-specific as long as it is exogenous to the counterfactuals, in the sense that $\tau_{i}$ is independent of $y_{it}(0)$ for each $i$. Our approach can then be applied to a staggered adoption setting where individuals choose the timing of the treatment at random. [^24] For a discussion, see Viviano and Bradic (2023). The basis functions ${b_{k}(t)}$ in Definition 1 are then functions of the time to adoption $t$ and the adoption period is set $\tau_{i}=0$. Additionally, it is possible to allow for treatment anticipation, as long as it is limited. In this case, one can simply modify the pre-treatment estimation window ${\cal T}_{i}$ in Definition 1 to include observations only up to the time $\tau-\delta_{i}$ at which it is still reasonable to assume that there was no treatment anticipation, that is, ${\cal T}_{i}\equiv\left\{\tau-\delta_{i}-R_{i}+1,\ldots,\tau-\delta_{i}\right\}$ (and $h$ is adjusted accordingly).
2.5 Balanced panel and pooled estimation
Assume that $R=R_{i}$ and $q=q_{i}$ are constant across $i$ and focus on polynomial basis functions in Definition 1. The first alternative way to obtain our proposed estimator, in this case, is to consider the cross-sectional averages $\overline{y}_{t}=\frac{1}{n}\sum_{i=1}^{n}y_{it}$ of the observed outcomes in time period $t$. Due to linearity of the forecasting procedure, we can rewrite:

### v4

@@ 2.2 Unbiased forecasts of counterfactuals
In this section, we characterize the class of DGPs for the counterfactuals. The need to discuss the DGP for counterfactuals arises because of the lack of an untreated group, which means that we must rely on forecasting counterfactuals from pre-treatment observations.
@@ 2.2.1 Stationary or stochastic trends DGPs
In this section, we consider a class of DGPs such that Assumption 1 is satisfied generally, namely, by any forecast that can be written as a weighted average of pre-treatment outcomes with weights summing to 1. The DGPs in this class express the counterfactual as the sum of potentially two unobserved stochastic components. This includes a variety of processes, such as stationary and non-stationary (unit root) ARMA processes with individual-specific parameters.
@@
where $y_{it}^{(1)}(0)$ is an unobserved mean-stationary process (i.e., $\mathbb{E}y_{it}^{(1)}(0)$ is constant over t) and $y_{it}^{(2)}(0)=y_{it-1}^{(2)}(0)+u_{it}(0)$ is an unobserved random walk process with $\mathbb{E}u_{it}(0)=0$ for all $t\geq 2$. Either/both components could be zero.
@@
Assumption 3 does not require both components to be present, which means that it accommodates stationarity as well as non-stationarity due to a stochastic trend. The user does not need to take a stance on the component(s). Our method is robust to both. When both components in Assumption 3 are present, the assumption is equivalent to the classical trend-cycle decomposition of macroeconomic time series with stochastic trends (e.g., Nelson and Plosser (1982); Watson (1986)).
@@
This class of DGPs is a plausible assumption for applications where either (1) the time series of pre-treatment outcomes does not display a trend; (2) there is a trend in pre-treatment outcomes that is plausibly stochastic (not deterministic); or (3) there is only one pre-treatment observation so a deterministic trend could never be modeled anyway. We consider processes with deterministic trends in pre-treatment outcomes in the next section.
A key insight of this paper is that a correctly specified parametric model (e.g., a specific ARIMA model) is not necessary to obtain unbiased forecasts of the counterfactuals. In fact, as the next result shows, any forecast expressed as a weighted average of pre-treatment observations satisfies the unbiasedness condition.
@@
Note that Theorem 1 shows how to obtain unbiased estimates of the individual (possibly heterogeneous) treatment effects. The result in (11) is stronger than the average unbiasedness required by Assumption 1. This means that under the assumptions of Theorem 1 one can not only obtain consistent and asymptotically normal estimates of the average treatment effects, but also unbiased estimates of the individual treatment effects.
There are many ways to obtain forecasts that are weighted averages of pre-treatment data. In this paper we focus on a general class of forecasts obtained via basis function regressions, such as polynomial time trends regressions.
@@
Consider a sequence of linearly independent functions $\{b_{k}(t)\}_{k=0}^{q_{i}}$, $q_{i}\in\left\{0,1,2,\ldots,\tau-1\right\}$, on the set $\mathcal{T}_{i}=\left\{\tau-R_{i}+1,\ldots,\tau\right\}$ with $R_{i}\in\left\{q_{i}+1,\ldots,\tau\right\}$, and such that $b_{0}(t)=1$ for all $t$. For example, polynomial time trends set $b_{k}(t)=t^{k}$, with $q_{i}$ the order of the polynomial. For each individual $i$, we forecast the counterfactual via individual-specific regressions of pre-treatment outcomes $\{y_{it}\}_{t\in\mathcal{T}_{i}}$ on the basis functions $\{b_{k}(t)\}_{k=0}^{q_{i}}$:
@@
This definition makes it clear that for any type of basis function the choice $q_{i}=0$ yields the sample mean of pre-treatment outcomes as the forecast. The following result shows that forecasts obtained via basis function regressions satisfy the weighted average requirement of Theorem 1.
@@
For known basis functions $\{b_{k}(t)\}_{k=0}^{q_{i}},\,q_{i}=0,1,\dots,\tau_{i}-1$ that are linearly independent on $\mathcal{T}_{i}$ with $b_{0}\left(t\right)=1$, the forecast in equation (12) satisfies equation (10).
Our framework connects to the literature on random coefficient panel models, particularly Chamberlain (1992), Graham and Powell (2012), and Arellano and Bonhomme (2012). Linear panel models with unit-specific intercepts and trend coefficients, similar to specifications in this literature, belong to the class of DGPs in Assumption 3 or Assumption 4. However, our goals differ: that literature studies identification of the distribution of heterogeneous coefficients, whereas we focus on forecasting untreated potential outcomes to estimate average treatment effects.
@@ 2.2.2 Deterministic trends DGPs
where $y_{it}^{(1)}(0)$ and $y_{it}^{(2)}(0)$ are as in Assumption 3 and $y_{it}^{(3)}(0)$ is a deterministic time trend $y_{it}^{(3)}(0)=\sum_{k=0}^{q_{0i}}c_{ik}^{(3)}{b}_{k}(t)$ with $c_{i}^{(3)}\in\mathbb{R}^{q_{0i}+1}$ and known basis functions $\{b_{k}(t)\}_{k=0}^{q_{0i}}$, $q_{0i}\in\left\{0,1,2,\ldots,\tau-1\right\}$.
@@
where $\widehat{y}_{i\tau+h}^{(q_{i},R_{i})}(0)$ is defined in (12). If $y_{it}^{(3)}(0)$ is not zero, it must be that $q_{i}\geq q_{0i}$.
@@
While the stochastic components of the counterfactual in Assumption 4 are unobserved and do not have to be present, if the deterministic time trend component is present, it must be a function of the same basis functions used to obtain the forecast. This implies that the stochastic component of the DGP does not need to be correctly specified, but, if a deterministic time trend component is present, it is correctly specified up to the order (as we discuss in the next remark). This is an unusual result from the perspective of forecasting, where one typically focuses on specifying both stochastic and deterministic parts of a model.
@@
The key requirement of Theorem 2 is that $q_{i}$ (the number of basis functions used in estimation) be greater than or equal to $q_{0i}$ (the true number of basis functions), when the DGP has a deterministic time trend component. Intuitively, this means that, in the presence of deterministic trends, choosing a too small number of basis functions runs the risk of delivering biased forecasts of the counterfactuals. We discuss the practical implications of these findings when discussing the choice of tuning parameters in Section 2.3.
Remark 6.
It may be possible to weaken the assumption that there is no forecastable common shock between $\tau$ and $\tau+h,h\geq 1$. To see this, let $\widetilde{y}_{it}\left(0\right)$ follow either Assumption 3 or 4 and let $W_{it}$ be an observed variable that is affected by a common shock $\gamma_{t}$ but not by the treatment. Suppose that
$W_{it}$ $=\theta_{i}^{W}\gamma_{t}1\left(t>\tau\right)+\epsilon_{it}.$ (16)
Denote by $\widehat{\gamma}_{t}$ an estimator of $\gamma_{t}$ obtained from post-treatment data on $W_{it},$ $t>\tau$ via, e.g., PCA. Then
where $\widehat{\widetilde{y}}_{i\tau+h}(0)$ is the counterfactual forecast obtained as before via basis function regression.
Leveraging auxiliary information to partial out the effect of unobserved aggregate shocks is conceptually related to White and Kennedy (2009); Freyaldenhoven et al. (2019); Brown et al. (2021), among others.
@@ 2.3 Choice of basis functions and tuning parameters
Our proposed method for forecasting counterfactuals in Definition 1 requires choosing: 1) the number of pre-treatment periods used for the estimation $R_{i}$; 2) the type of basis functions $b_{k}(t)$; and 3) the number of basis functions $q_{i}$. We discuss how these choices affect inference and offer some practical recommendations for empirical researchers.
The choice of estimation window, $R_{i}$, involves a trade-off. In our class of DGPs, a larger $R_{i}$ generally yields an estimator with smaller variance, but a shorter $R_{i}$ can guard against violations of our assumptions stemming from parameter instability in pre-treatment data. While plotting pre-treatment time series offers informal guidance on such instability, applying our method to long-T settings reveals a heightened sensitivity to the chosen estimation window. This is because the increased likelihood of structural changes and parameter instability over extended periods becomes a critical concern. This challenge resonates with discussions in the regression discontinuity literature regarding bandwidth selection (e.g., Gelman and Imbens 2019).
@@
Regarding the choice of basis functions, Theorems 1 and 2 imply that this choice only matters for unbiasedness when the DGP has a deterministic time trend, in which case the basis functions need to be correctly specified to ensure unbiasedness (up to the order). When the DGP is mean stationary or has a stochastic trend, the choice of basis functions does not matter for unbiasedness. Basis functions may be chosen based on the time series properties of pre-treatment outcomes. Polynomial time trends seem to be a natural choice of basis functions for DGPs with deterministic trends. [^10] In applications using difference-in-differences (DiD) methods it is typical to assume the presence of time trends (mostly linear) that are common between control and treatment groups. Our results make it clear that a linear time trend can only be dealt with by either using an untreated group or, when an untreated group is not available, by using (at least two) pre-treatment time periods to model the trend (which leads to our polynomial regression). Other basis functions could be used, e.g., periodicity could be captured by Fourier basis functions. Our practical recommendation, and what we focus on henceforth, is to consider polynomial basis functions by letting $b_{k}(t)=t^{k}$ in Definition 1.111111An advantage of polynomial basis functions is that it is in principle possible to relax the assumption that the basis functions are correctly specified by instead assuming that $y_{it}^{(3)}(0)$ in Assumption 4 is a continuous (but otherwise unspecified) function of time. Since time in our setting is defined on a compact interval and the deterministic trend is a continuous function, $y_{it}^{(3)}(0)$ can be approximated arbitrarily well by a polynomial in time. In fact, by the Weierstrass Approximation Theorem, the approximation error approaches zero as the order of the polynomial goes to infinity. Under additional smoothness assumptions on the deterministic trend, an approximation theorem could then be used (e.g. the Polynomial Approximation Error Theorem) to derive a bound on the approximation error of $y_{it}^{(3)}(0)$ by the polynomial regression. The forecast of $y_{i\tau+h}(0)$ is biased, but we conjecture that it may be possible to do bias correction given an expression for the bias obtained via the approximation theorem.
Regarding the choice of the order of the basis functions $q_{i}$ used for the estimation, this only matters for unbiasedness if the DGP has a deterministic trend component. For DGPs without such a trend, any $q_{i}$ ensures unbiasedness, with $q_{i}=0$ offering the lowest variance; however, larger $q_{i}$ can mitigate bias from non-stationary initial conditions. If a deterministic trend exists, $q_{i}$ cannot be smaller than the unknown true order, forcing a trade-off where larger $q_{i}$ for unbiasedness risks higher variance. This choice is inherently constrained by the number of pre-treatment periods ($T$): larger $T$ allows for more flexible polynomial forecasts and extensive placebo testing, yet estimation consumes periods from this budget; conversely, smaller $T$ drastically limits both model flexibility and the scope for placebo tests.
Standard cross-validation methods typically target predictive accuracy, not the unbiasedness crucial for our context. While bias-targeting cross-validation is a promising avenue for selecting $q_{i}$, its implementation faces challenges in our short-T panel context. It inherently requires long pre-treatment series for data splitting, which is often not feasible. Moreover, even with long-T data, this can heighten concerns about parameter instability (as stationarity becomes less defensible over extended periods). Therefore, our practical recommendation is to report results for a small range of $q_{i}$ values (e.g., $0,1,2,3$), with pre-treatment time series plots providing informal guidance.
@@ 2.4 Individual treatment timing and limited anticipation
The treatment timing $\tau_{i}$ can be individual-specific. While staggered adoption designs typically allow for identification via comparisons with not-yet-treated units, such strategies are inherently limited to the window before the last unit becomes treated. In settings where all units eventually adopt the treatment, such as the universal rollout of a national policy, standard event-study designs cannot recover treatment effects for the periods following full adoption, as no valid comparison group remains.
Our approach overcomes this limitation. Because FAT constructs counterfactuals using only the treated unit’s own pre-treatment history, it does not rely on the existence of a contemporaneous control group. Consequently, it allows researchers to estimate treatment effects over longer horizons, extending into periods after all units have been treated.
We require that the treatment timing be exogenous with respect to the counterfactual outcome path $y_{it}(0)$, conditional on the unit’s history used for forecasting. The basis functions $\{b_{k}(t)\}$ in Definition 1 are then functions of the time to adoption $t-\tau_{i}$. Additionally, it is possible to allow for treatment anticipation, as long as it is limited. In this case, one can simply modify the pre-treatment estimation window ${\cal T}_{i}$ in Definition 1 to include observations only up to the time $\tau_{i}-\delta_{i}$ at which it is still reasonable to assume that there was no treatment anticipation, that is, ${\cal T}_{i}\equiv\left\{\tau_{i}-\delta_{i}-R_{i}+1,\ldots,\tau_{i}-\delta_{i}\right\}$ (and $h$ is adjusted accordingly).
2.5 Balanced panels and pooled estimation
For balanced panels, our proposed estimator is algebraically equivalent to two simpler pooled estimation strategies. These alternative approaches are also suitable for repeated cross-section data or when cohort of birth plays the role of time. However, in unbalanced panels, these simpler strategies may yield inconsistent estimates for ATT. Therefore, our exposition primarily focuses on individual-level estimators, given their ability to accommodate unbalanced panels; these same pooled estimators, however, are how the FAT is obtained in balanced panels.
Assume that $R=R_{i}$ and $q=q_{i}$ are constant across $i$ and focus on polynomial basis functions in Definition 1. The first alternative way to obtain our proposed estimator is to consider the cross-sectional averages $\overline{y}_{t}=\frac{1}{n}\sum_{i=1}^{n}y_{it}$ of the observed outcomes in time period $t$. Due to linearity of the forecasting procedure, we can rewrite: